# InsurTech innovation using natural language processing

### v2

@@ 4 NLP in Action: Use-cases Leveraging InsurTech Datasets
[108] word2vec/GloVe Loss Modeling Data Enrichment
[112] GloVe Loss Modeling Data Enrichment
[76] TF-IDF $\mathcal{N}$- grams/GloVe Loss Modeling Data Enrichment
[147] $\mathcal{N}$-gram Information Retrieval Data Enrichment
[81] word2vec - Data Enrichment
[113] word2vec Loss Modeling Data Enrichment
[145] BERT Loss Modeling Data Enrichment
[129] fastText Loss Modeling Data Enrichment
[144] BERT Loss Modeling Data Enrichment
[137] DistilBERT Loss Modeling Data Enrichment
[79] $\mathcal{N}$-gram Sentiment Analysis Data Enrichment
[100] L-M sentiment dictionary Text Classification Data Enrichment
@@
[110] TF-IDF Customer Service Standalone Application
[77] LLM Document Summarization Standalone Application
[148] Sentence-BERT Information Retrieval Standalone Application
@@
[86] BERT/BGE/LLaMA3 Text Classification Standalone Application
[109] LLM Sentiment Analysis Standalone Application
[105] LLM Fraud Detection Standalone Application
[143] LLM Claim Management Standalone Application
@@ 4.1 InsurTech Data
InsurTech, much like FinTech, has emerged as a key catalyst in transforming the conventional insurance sector into a modern, data-driven industry in the era of big data and AI. This transformation introduces a diversity of data sources and formats, among which textual data has become one of the most prominent and extensively studied in the current AI landscape. A vast amount of textual data can be scraped from the internet, including information that businesses provide about their products and services on their websites, as well as customer reviews and feedback shared on social media and consumer rating platforms. While this data is relatively easy to obtain, extracting meaningful insights from it remains a challenge. However, advancements in NLP techniques offer powerful tools to analyze this unstructured text effectively. These methods allow businesses to gain a deeper understanding of their operations and customers’ perceptions of their products and services. For insurers, leveraging these insights can enhance risk management by identifying potential vulnerabilities, improving customer segmentation, and refining underwriting and claims processes. In this work, we extend the real-world InsurTech dataset introduced in prior research [122], by incorporating additional textual data to support our feature de-biasing and industry classification use cases. We obtained the raw data from our InsurTech partner and applied a series of data preprocessing steps, as discussed in Section 3, to prepare the data for downstream predictive modeling. Since our study incorporates only a limited set of numerical variables from prior work, and the majority of our NLP use cases rely exclusively on newly introduced textual data, we omit unnecessary details of irrelevant numerical variables and instead discuss the adopted numerical features within the context of the corresponding use cases. To demonstrate the value of textual data in enhancing insurance operations, we specifically incorporate business reviews scraped from crowd-sourced platforms (e.g., Google Maps and Yelp user reviews) as well as descriptive texts provided by insured businesses on their websites. Regarding the textual components, since the InsurTech data is proprietary and the majority of the textual information contains easily identifiable details about the underlying businesses, we are unable to present the raw text directly. Nevertheless, in our use cases, we provide selected examples that illustrate the data while ensuring that no sensitive information is disclosed.
@@ 4.2.1 Data Enrichment Using Text
As discussed extensively in Section 3, converting textual data into numerical representations is a well-established tradition that continues to evolve. In actuarial science, numerous efforts have been made to leverage textual data to enhance predictive modeling in insurance. For instance, [108] introduce a framework that incorporates textual data into insurance claims modeling and examine its applications in claims management processes. Building on this foundation, [112] and [113] present a series of studies that utilized embedding techniques to enhance pricing models. Similarly, [147] demonstrate the application of NLP techniques in insurance by analyzing accident narratives from police reports to classify risk profiles and fine-tune policy premiums. [76] propose a data-driven approach that automatically extracts actuarial risk factors from a large corpus of accident descriptions, further emphasizing the potential of NLP in refining insurance models. Expanding on these efforts, [145] and [144] employ BERT embeddings to enhance the frequency and severity prediction of truck warranty claims, underscoring the transformative impact of NLP in improving predictive models for actuarial tasks. [129] extend the use of text-based claim predictions to imbalanced datasets by developing a framework that leverages neural networks to automatically identify extremely severe claims. [137] embed car accident descriptions into condensed numerical vectors using DistilBERT, which are then incorporated as features to enrich the classical Generalized Linear Model (GLM) framework for insurance pricing. [81] provide an overview of textual embedding techniques within a unified neural network–based data representation framework. [79] propose RMQuality, a textual-analysis framework designed to quantify firms’ climate change risk management quality, thereby assisting firms in identifying and mitigating climate risks. Moreover, [100] identify firms’ risk culture through engineered textual-informed features and examine its correlation with regulatory practices. Collectively, these works highlight the growing role of NLP in actuarial science, offering innovative approaches to improve predictive accuracy, risk assessment, and decision-making in the insurance industry.
@@ 4.2.2 De-Biasing Features Through Sentiment Analysis
In our study, we identified a potential bias in the review star rating system, which could compromise the reliability of supervised learning models. Since different individuals may assign different star values despite expressing similar sentiments, using these ratings as labels for training an ML model could introduce bias. To mitigate this issue, we adopt the Lexicon-Based Approach for our purpose. The lexicon-based method typically begins with a seed set of words that carry known sentiment meanings. These seed words are then expanded using large synonym dictionaries such as WordNet [^3] https://wordnet.princeton.edu/, allowing additional words with similar sentiment features to be incorporated into the lexicon [141]. Another approach to constructing a sentiment lexicon leverages statistical and semantic techniques. By analyzing semantic relationships, word similarity, and Pointwise Mutual Information (PMI), it is possible to assign sentiment scores to words in a corpus, thereby building a lexicon with sentiment polarity. Research has shown that using word similarity, lexical association, and, most importantly, distributional similarity, sentiment polarity scores can be systematically assigned to words in a corpus [127]. By applying a lexicon-based sentiment analysis approach, we can generate a more objective sentiment scoring system based on textual content, reducing reliance on the potentially biased star rating system. Given the word, we can retrieve its corresponding sentiment polarity score, which typically ranges between -1 and 1. To calculate the sentiment of a sentence, we average the sentiment score of all words in the sentence using TextBlob [^4] https://textblob.readthedocs.io/en/dev/ package. It derives sentiment scores for each word in our business reviews from SentiWordNet [75], a lexical dictionary, and calculates the overall polarity score as the average sentiment score of all words.
@@ 4.2.3 Transforming High-Cardinality Categorical Features into Numerical Embeddings
In real-world predictive modeling, it is common to encounter high-cardinality categorical features where the number of distinct categories is large, while the dataset itself is relatively small. This inequality often leads to the curse of dimensionality, a phenomenon that is particularly pronounced in academic settings where the available data is significantly smaller compared to industry datasets. Traditionally, researchers have addressed this challenge by applying regularization penalties to control model complexity. However, recent advances in NLP offer an alternative approach. [132] demonstrate the use of categorical embeddings to transform categorical features into numerical vectors within the non-life insurance domain, enabling more effective handling of high-cardinality features. Building on this, [122] explore the feasibility of incorporating InsurTech features to enhance business insurance pricing. Specifically, prior work aims to reduce the dimensionality of business categorization by employing a combination of text embedding and clustering techniques.
@@
Our analysis reveals that text embedding and clustering techniques have been proven inadequate for effective feature compression. The InsurTech dataset includes raw business classifications extracted from various websites. These raw classifications are highly complex, consisting of 13,287 unique categories with unprocessed, messy text labels, making them difficult to use effectively. In our prior work [122], to prevent a drastic increase in dimensionality, these unique business categories have been first embedded into 100-dimensional vectors using word2vec, capturing their static semantic meanings. The embeddings have then been clustered into 24 groups via k-means clustering, aiming to aggregate semantically similar categories. Table 2 presents ten such business categories and corresponding category clusters. While this approach has provided a way to encode highly complex categorical variables into more manageable, low-dimensional categories, it introduces two significant drawbacks. First, some categories, for example, 1- to 5-star hotels, have been grouped into the same cluster (cluster 5) in previous experiments. While these categories share semantic similarities, they may present very different underlying risks. This illustrates that the clustering approach preserves semantic similarity but loses granularity in risk differentiation. Second, static embeddings often fail to capture subtle semantic variations in the category titles, introducing information loss. For instance, exchanging the words “Repair” and “Service” does not alter the clustering results (cluster 10). However, introducing slight modifications, such as introducing “Installation” or “Replacement,” pluralizing “Repair” to “Repairs,” or even using “A/C” instead of “AC,” can significantly change the resulting cluster assignment. This sensitivity to superficial textual differences is undesirable, as it compromises the robustness and interpretability of the classification. This limitation is not unique to the approaches discussed here; the literature referenced in Section 4.2.1 may also suffer from similar drawbacks. Therefore, instead of collapsing high-dimensional categorical features into lower-dimensional categorical clusters using static embedding, a more effective approach is to transform them into low-dimensional numerical features using context-aware embedding techniques.
@@ 5 Star(s) Hotel 5
To further validate the effectiveness of categorical embeddings, we revisit the empirical experiments presented in [122], replacing the 24 cluster-based categorical features with 24 numerical features derived through direct embedding of business categories. Unlike earlier techniques that rely on static semantic representations (e.g., word2vec or GloVe), modern embedding models are trained on vast amounts of textual data, enabling them to capture deeper contextual and semantic nuances. In this study, we adopt the BGE-M3 embedding model introduced by [87], a state-of-the-art multilingual and multifunctional embedding model that produces 1,024-dimensional numerical representations for each business category. As general-purpose LLMs exhibit remarkable transfer learning capabilities, modern NN–based embedding models offer a powerful approach to leverage textual data. Unlike conventional static embeddings, these models generate more effective and informative numerical representations, thereby improving predictive performance in downstream tasks.
@@
However, due to the high dimensionality of these embeddings, direct use in downstream predictive modeling may lead to overfitting or inefficiencies. To address this, we implement the PCA-based dimensionality reduction technique proposed by [126], selecting the top 24 principal components that explain the greatest variance. This enables us to construct 24 context-aware numerical features that preserve the core semantic information while maintaining compatibility with the experimental setup in [122]. As an illustrative example, Figure 5 displays the first two principal components of the embeddings for ten category titles (the same as those in Table 4), since the full 24-dimensional vectors are not easily visualizable. It is important to note that because only the first two components are retained and the vector lengths are adjusted for visualization purposes, the figure does not perfectly reflect the actual embeddings used in later empirical experiments. Nevertheless, the plot highlights that category titles with similar semantic meanings yield closer embeddings. For instance, hotel-related titles generally appear in the opposite direction of AC system–related titles. Unlike the static embedding coupled with a clustering approach, the NN-based embeddings capture subtle semantic variations. The lower-star hotels (1- and 2-star) are positioned toward the left of the two-dimensional space, while higher-star hotels (4- and 5-star) shift toward the right. Interestingly, “A/C” and “AC” exhibit greater divergence than expected, whereas word changes such as “Installation,” “Replacement,” and “Service,” which all refer to similar electrical appliance services, produce only minor differences. These results demonstrate that modern neural embedding models, trained on vast textual corpora, can effectively transform high-dimensional categorical variables into practical numerical embeddings while addressing the two key challenges identified in prior approaches.
@@
We then retrain the model with the numerical embedded vectors as numerical features. To evaluate model performance and ensure comparability with prior studies, we adopt four commonly used metrics: Gini index, Percentage Error (PE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) (detailed in Appendix C). Specifically, the Gini index assesses the discrimination in prediction order, PE captures portfolio predictive uncertainty, and RMSE and MAE quantify individual-level predictive accuracy, with RMSE placing greater emphasis on tail distributions. Together, these metrics provide a comprehensive evaluation of model predictive capabilities. As summarized in Table 5, the insurance in-house model refers to the data provider’s proprietary pricing model, which serves as the baseline. LightGBM with category clustering represents the original approach that applies clustering as a dimensionality reduction technique, both reported in the BG section of Table 4 in [122]. By contrast, LightGBM with category embedding introduces the new numerical embedding method, where high-dimensional categorical features are transformed into low-dimensional, context-aware numerical embeddings, leading to improved model performance. It should be noted that the BG coverage includes 561 features, of which the 24 category clustering/embedding features constitute only a small fraction of the overall feature space. The rest of the numerical 537 features, such as additional risk characteristics and industry indicators, also contribute to interpreting the business category, thereby reducing the relative influence of our engineered features. Consequently, the performance gain of category embedding is expected to be moderate at best. Moreover, the complexity of real-world InsurTech datasets further limits the potential for substantial metric improvements. Although we retain only 24 numerical features for consistency, further increasing the number of retained components may yield additional gains, leaving space for further significant improvement. These findings reinforce the value of NLP-driven categorical embedding techniques for transforming high-cardinality categorical variables into compact, informative numerical features, ultimately improving predictive modeling in insurance by better capturing underlying semantic structures.

### v3

@@ 4 NLP in Action: Use-cases Leveraging InsurTech Datasets
[Lee2020] word2vec/GloVe Loss Modeling Data Enrichment
[manski_extracting_2021] GloVe Loss Modeling Data Enrichment
[baillargeon2021mining] TF-IDF $\mathcal{N}$- grams/GloVe Loss Modeling Data Enrichment
[zappa_text_2021] $\mathcal{N}$-gram Information Retrieval Data Enrichment
[manski_loss_2022] word2vec Loss Modeling Data Enrichment
[xu_bert-based_2022] BERT Loss Modeling Data Enrichment
[sabban2022automatic] fastText Loss Modeling Data Enrichment
[xu2023framework] BERT Loss Modeling Data Enrichment
[troxler2024actuarial] DistilBERT Loss Modeling Data Enrichment
[berry-stolzleInsurersClimateChange2024] $\mathcal{N}$-gram Sentiment Analysis Data Enrichment
[guptaRegulatingRiskCulture2025] L-M sentiment dictionary Text Classification Data Enrichment
@@
[Liao2020] TF-IDF Customer Service Standalone Application
[Zhang11112024] Sentence-BERT Information Retrieval Standalone Application
@@
[cao2024assessing] BERT/BGE/LLaMA3 Text Classification Standalone Application
[liTextualAnalysisInsurance2025] LLM Sentiment Analysis Standalone Application
@@ 4.1 InsurTech Data
InsurTech, much like FinTech, has emerged as a key catalyst in transforming the conventional insurance sector into a modern, data-driven industry in the era of big data and AI. This transformation introduces a diversity of data sources and formats, among which textual data has become one of the most prominent and extensively studied in the current AI landscape. A vast amount of textual data can be scraped from the internet, including information that businesses provide about their products and services on their websites, as well as customer reviews and feedback shared on social media and consumer rating platforms. While this data is relatively easy to obtain, extracting meaningful insights from it remains a challenge. However, advancements in NLP techniques offer powerful tools to analyze this unstructured text effectively. These methods allow businesses to gain a deeper understanding of their operations and customers’ perceptions of their products and services. For insurers, leveraging these insights can enhance risk management by identifying potential vulnerabilities, improving customer segmentation, and refining underwriting and claims processes. In this work, we extend the real-world InsurTech dataset introduced in prior research [Quan15102024], by incorporating additional textual data to support our feature de-biasing and industry classification use cases. We obtained the raw data from our InsurTech partner and applied a series of data preprocessing steps, as discussed in Section 3, to prepare the data for downstream predictive modeling. Since our study incorporates only a limited set of numerical variables from prior work, and the majority of our NLP use cases rely exclusively on newly introduced textual data, we omit unnecessary details of irrelevant numerical variables and instead discuss the adopted numerical features within the context of the corresponding use cases. To demonstrate the value of textual data in enhancing insurance operations, we specifically incorporate business reviews scraped from crowd-sourced platforms (e.g., Google Maps and Yelp user reviews) as well as descriptive texts provided by insured businesses on their websites. Regarding the textual components, since the InsurTech data is proprietary and the majority of the textual information contains easily identifiable details about the underlying businesses, we are unable to present the raw text directly. Nevertheless, in our use cases, we provide selected examples that illustrate the data while ensuring that no sensitive information is disclosed.
@@ 4.2.1 Data Enrichment Using Text
As discussed extensively in Section 3, converting textual data into numerical representations is a well-established tradition that continues to evolve. In actuarial science, numerous efforts have been made to leverage textual data to enhance predictive modeling in insurance. For instance, [Lee2020] introduce a framework that incorporates textual data into insurance claims modeling and examine its applications in claims management processes. Building on this foundation, [manski_extracting_2021] and [manski_loss_2022] present a series of studies that utilized embedding techniques to enhance pricing models. Similarly, [zappa_text_2021] demonstrate the application of NLP techniques in insurance by analyzing accident narratives from police reports to classify risk profiles and fine-tune policy premiums. [baillargeon2021mining] propose a data-driven approach that automatically extracts actuarial risk factors from a large corpus of accident descriptions, further emphasizing the potential of NLP in refining insurance models. Expanding on these efforts, [xu_bert-based_2022] and [xu2023framework] employ BERT embeddings to enhance the frequency and severity prediction of truck warranty claims, underscoring the transformative impact of NLP in improving predictive models for actuarial tasks. [sabban2022automatic] extend the use of text-based claim predictions to imbalanced datasets by developing a framework that leverages neural networks to automatically identify extremely severe claims. [troxler2024actuarial] embed car accident descriptions into condensed numerical vectors using DistilBERT, which are then incorporated as features to enrich the classical Generalized Linear Model (GLM) framework for insurance pricing. [berry-stolzleInsurersClimateChange2024] propose RMQuality, a textual-analysis framework designed to quantify firms’ climate change risk management quality, thereby assisting firms in identifying and mitigating climate risks. Moreover, [guptaRegulatingRiskCulture2025] identify firms’ risk culture through engineered textual-informed features and examine its correlation with regulatory practices. Collectively, these works highlight the growing role of NLP in actuarial science, offering innovative approaches to improve predictive accuracy, risk assessment, and decision-making in the insurance industry.
@@ 4.2.2 De-Biasing Features Through Sentiment Analysis
In our study, we identified a potential bias in the review star rating system, which could compromise the reliability of supervised learning models. Since different individuals may assign different star values despite expressing similar sentiments, using these ratings as labels for training an ML model could introduce bias. To mitigate this issue, we adopt the Lexicon-Based Approach for our purpose. The lexicon-based method typically begins with a seed set of words that carry known sentiment meanings. These seed words are then expanded using large synonym dictionaries such as WordNet [^3] https://wordnet.princeton.edu/, allowing additional words with similar sentiment features to be incorporated into the lexicon [Whitelaw2005]. Another approach to constructing a sentiment lexicon leverages statistical and semantic techniques. By analyzing semantic relationships, word similarity, and Pointwise Mutual Information (PMI), it is possible to assign sentiment scores to words in a corpus, thereby building a lexicon with sentiment polarity. Research has shown that using word similarity, lexical association, and, most importantly, distributional similarity, sentiment polarity scores can be systematically assigned to words in a corpus [Read2009]. By applying a lexicon-based sentiment analysis approach, we can generate a more objective sentiment scoring system based on textual content, reducing reliance on the potentially biased star rating system. Given the word, we can retrieve its corresponding sentiment polarity score, which typically ranges between -1 and 1. To calculate the sentiment of a sentence, we average the sentiment score of all words in the sentence using TextBlob [^4] https://textblob.readthedocs.io/en/dev/ package. It derives sentiment scores for each word in our business reviews from SentiWordNet [baccianellaSentiWordNet30Enhanced2010], a lexical dictionary, and calculates the overall polarity score as the average sentiment score of all words.
@@ 4.2.3 Transforming High-Cardinality Categorical Features into Numerical Embeddings
In real-world predictive modeling, it is common to encounter high-cardinality categorical features where the number of distinct categories is large, while the dataset itself is relatively small. This inequality often leads to the curse of dimensionality, a phenomenon that is particularly pronounced in academic settings where the available data is significantly smaller compared to industry datasets. Traditionally, researchers have addressed this challenge by applying regularization penalties to control model complexity. However, recent advances in NLP offer an alternative approach. [shiNonLifeInsuranceRisk2023] demonstrate the use of categorical embeddings to transform categorical features into numerical vectors within the non-life insurance domain, enabling more effective handling of high-cardinality features. Building on this, [Quan15102024] explore the feasibility of incorporating InsurTech features to enhance business insurance pricing. Specifically, prior work aims to reduce the dimensionality of business categorization by employing a combination of text embedding and clustering techniques.
@@
Our analysis reveals that text embedding and clustering techniques have been proven inadequate for effective feature compression. The InsurTech dataset includes raw business classifications extracted from various websites. These raw classifications are highly complex, consisting of 13,287 unique categories with unprocessed, messy text labels, making them difficult to use effectively. In our prior work [Quan15102024], to prevent a drastic increase in dimensionality, these unique business categories have been first embedded into 100-dimensional vectors using word2vec, capturing their static semantic meanings. The embeddings have then been clustered into 24 groups via k-means clustering, aiming to aggregate semantically similar categories. Table 2 presents ten such business categories and corresponding category clusters. While this approach has provided a way to encode highly complex categorical variables into more manageable, low-dimensional categories, it introduces two significant drawbacks. First, some categories, for example, 1- to 5-star hotels, have been grouped into the same cluster (cluster 5) in previous experiments. While these categories share semantic similarities, they may present very different underlying risks. This illustrates that the clustering approach preserves semantic similarity but loses granularity in risk differentiation. Second, static embeddings often fail to capture subtle semantic variations in the category titles, introducing information loss. For instance, exchanging the words “Repair” and “Service” does not alter the clustering results (cluster 10). However, introducing slight modifications, such as introducing “Installation” or “Replacement,” pluralizing “Repair” to “Repairs,” or even using “A/C” instead of “AC,” can significantly change the resulting cluster assignment. This sensitivity to superficial textual differences is undesirable, as it compromises the robustness and interpretability of the classification. This limitation is not unique to the approaches discussed here; the literature referenced in Section 4.2.1 may also suffer from similar drawbacks. Therefore, instead of collapsing high-dimensional categorical features into lower-dimensional categorical clusters using static embedding, a more effective approach is to transform them into low-dimensional numerical features using context-aware embedding techniques.
@@ 5 Star(s) Hotel 5
To further validate the effectiveness of categorical embeddings, we revisit the empirical experiments presented in [Quan15102024], replacing the 24 cluster-based categorical features with 24 numerical features derived through direct embedding of business categories. Unlike earlier techniques that rely on static semantic representations (e.g., word2vec or GloVe), modern embedding models are trained on vast amounts of textual data, enabling them to capture deeper contextual and semantic nuances. In this study, we adopt the BGE-M3 embedding model introduced by [chenM3EmbeddingMultiLingualityMultiFunctionality2024], a state-of-the-art multilingual and multifunctional embedding model that produces 1,024-dimensional numerical representations for each business category. As general-purpose LLMs exhibit remarkable transfer learning capabilities, modern NN–based embedding models offer a powerful approach to leverage textual data. Unlike conventional static embeddings, these models generate more effective and informative numerical representations, thereby improving predictive performance in downstream tasks.
@@
However, due to the high dimensionality of these embeddings, direct use in downstream predictive modeling may lead to overfitting or inefficiencies. To address this, we implement the PCA-based dimensionality reduction technique proposed by [raunakEffectiveDimensionalityReduction2019], selecting the top 24 principal components that explain the greatest variance. This enables us to construct 24 context-aware numerical features that preserve the core semantic information while maintaining compatibility with the experimental setup in [Quan15102024]. As an illustrative example, Figure 5 displays the first two principal components of the embeddings for ten category titles (the same as those in Table 4), since the full 24-dimensional vectors are not easily visualizable. It is important to note that because only the first two components are retained and the vector lengths are adjusted for visualization purposes, the figure does not perfectly reflect the actual embeddings used in later empirical experiments. Nevertheless, the plot highlights that category titles with similar semantic meanings yield closer embeddings. For instance, hotel-related titles generally appear in the opposite direction of AC system–related titles. Unlike the static embedding coupled with a clustering approach, the NN-based embeddings capture subtle semantic variations. The lower-star hotels (1- and 2-star) are positioned toward the left of the two-dimensional space, while higher-star hotels (4- and 5-star) shift toward the right. Interestingly, “A/C” and “AC” exhibit greater divergence than expected, whereas word changes such as “Installation,” “Replacement,” and “Service,” which all refer to similar electrical appliance services, produce only minor differences. These results demonstrate that modern neural embedding models, trained on vast textual corpora, can effectively transform high-dimensional categorical variables into practical numerical embeddings while addressing the two key challenges identified in prior approaches.
@@
We then retrain the model with the numerical embedded vectors as numerical features. To evaluate model performance and ensure comparability with prior studies, we adopt four commonly used metrics: Gini index, Percentage Error (PE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) (detailed in Appendix C). Specifically, the Gini index assesses the discrimination in prediction order, PE captures portfolio predictive uncertainty, and RMSE and MAE quantify individual-level predictive accuracy, with RMSE placing greater emphasis on tail distributions. Together, these metrics provide a comprehensive evaluation of model predictive capabilities. As summarized in Table 5, the insurance in-house model refers to the data provider’s proprietary pricing model, which serves as the baseline. LightGBM with category clustering represents the original approach that applies clustering as a dimensionality reduction technique, both reported in the BG section of Table 4 in [Quan15102024]. By contrast, LightGBM with category embedding introduces the new numerical embedding method, where high-dimensional categorical features are transformed into low-dimensional, context-aware numerical embeddings, leading to improved model performance. It should be noted that the BG coverage includes 561 features, of which the 24 category clustering/embedding features constitute only a small fraction of the overall feature space. The rest of the numerical 537 features, such as additional risk characteristics and industry indicators, also contribute to interpreting the business category, thereby reducing the relative influence of our engineered features. Consequently, the performance gain of category embedding is expected to be moderate at best. Moreover, the complexity of real-world InsurTech datasets further limits the potential for substantial metric improvements. Although we retain only 24 numerical features for consistency, further increasing the number of retained components may yield additional gains, leaving space for further significant improvement. These findings reinforce the value of NLP-driven categorical embedding techniques for transforming high-cardinality categorical variables into compact, informative numerical features, ultimately improving predictive modeling in insurance by better capturing underlying semantic structures.